40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Mar 3 04:41:45 2015, Jenne, Update, LSC, Transitioned DARM to AS55Q, some other work 
    Reply  Tue Mar 3 19:19:54 2015, Jenne, Update, LSC, Transitioned DARM to AS55Q, some other work ALSfool_PRFPMI_2Mar2015.pdfALSfool_kindaEngaged_2March2015_noFrames.png
       Reply  Wed Mar 4 00:50:36 2015, rana, Bureaucracy, Treasure, Transitioned DARM to AS55Q, some other work 
Message ID: 11095     Entry time: Tue Mar 3 19:19:54 2015     In reply to: 11090     Reply to this: 11096
Author: Jenne 
Type: Update 
Category: LSC 
Subject: Transitioned DARM to AS55Q, some other work 

[Jenne, EricQ]

A slightly more coherent elog for last night's work.


All night, we've been using REFL165 to hold the PRMI.  It's working very nicely.  To help it catch lock, I've set the gain in the PRCL filter bank high, and then the *0.6 filter triggers on.  The carm_cm_up script now will lock the PRMI on REFL165. 

We had to reset the REFL165 phase after we acquired lock - it was -91, but now is -48.  I'm not sure why it changed so significantly from the PRMI-only config to the PRFPMI config.


We measured the ALS fool cancellation with the arms held off resonance, at arm powers of a few.  Although, they were moving around a lot, but the measurement stayed nice and smooth.  Anyhow, we get almost as good of cancellation as we saw with the single arm (after we made sure that both phase trackers had the same UGF):

ALSfool_PRFPMI_2Mar2015.pdf

We were able to partly engage it one time, but we lost lock at some point.  Since the frame builder / daqd decided that that would be just the *perfect* time to crash and restart, we don't have any frame data for this time.  We can see up to a few seconds before the lockloss, while we were ramping up the RF PD loop gain though, and MICH was hitting the rails.  I'm not sure if that's what caused the lockloss, but it probably didn't help.

The ALS fool gain was 22, and we were using FMs 4, 6 (the pendulum and Rana's "comp1"), the same filters that were used for the single arm case.  The LSC-MC filter bank gain lost lock when we got to about 5.6 (we were taking +3dB steps).

We were using REFL55I/(TRX + TRY) as our CARM RF error signal.  We were using REFL55 rather than REFL11 because we were worried that REFL11 didn't look good - maybe it was saturating or something.  To be looked into.

Here's the striptool that was running at the time, since we don't have frame data:


At this point, since we weren't sure what the final gain should be for the RF CARM signal, and we could sit at nice high arm powers (arm powers of 30ish correspond to CARM offsets of about 50pm), we decided to try just a straight jump over to the RF signals. 

The first time around, we jumped CARM to (-0.2)*REFL55/(TRX+TRY), but we only stayed lock for 1 or 2 seconds.  That was around 1:55am.

We decided that perhaps it would be good to get DARM moved over first, since it has a much wider linewidth, so the rest of the trials for the night were transitioning DARM over to (0.0006)*AS55Q/(TRX + TRY).  AS55 was saturating, so we reduced its analog gain from 18dB to 9dB and re-ran the LSC offsets.  The MICH noise was pretty high when we were at low CARM offset, although we noticed it more when DARM was on AS55.  In particular, there is some peak just below 40Hz that is causing a whole comb of harmonics, and dominating the MICH, PRC and DARM spectra.  I will try to get a snapshot of that tonight - I don't think we saved any spectra from last night.  Turning off DARM's FM3 boost helped lower the MICH noise, so we think that the problem is significant coupling between the two degrees of freedom. 

After the first one or two tries of getting DARM to AS55, we started engaging the arm ASC loops - they helped the dark port contrast considerably.  The POP spot still moves around, but the dark port gets much darker, and is more symmetric with the ASC on.

Attachment 1: ALSfool_PRFPMI_2Mar2015.pdf  6 kB  | Hide | Hide all | Show all
ALSfool_PRFPMI_2Mar2015.pdf
Attachment 2: ALSfool_kindaEngaged_2March2015_noFrames.png  42 kB  | Show | Hide all | Show all
ELOG V3.1.3-